The IMF projects that China’s real GDP per person will be 57,163 yuan in 2017 and 60,334 yuan in 2018 and that India’s real GDP per person will be 98,028 rupees in 2017 and 104,191 rupees in 2018. By maintaining their current growth rates, which country will be first to double its standard of living and when will that happen?



1
Expert's answer
2021-11-09T10:45:26-0500

CHINA

In 2017, the real GDP per person was 57,163 Yuan.

In 2018, the real GDP per person was 60,334 Yuan.

Calculate the real GDP per person growth rate -

Growth rate = [(Real GDP per person in 2018 - Real GDP per person in 2017) "\div" Real GDP per person in 2017] "\times" 100

"\frac{60,334 - 57,163}{57,163} \times 100 = growth \space rate"

5.55 % growth rate

China's real GDP per capita is growing at a rate of 5.55 %

Calculate the number of years it will take to double real GDP per person -

Number of years = 70"\div" Real GDP per person growth rate "= \frac{70}{5.55} = 12.61\space years"


In around 13 years, China would be able to double its real GDP per person.

It means that the China would be able to double its real GDP per person by 2031.


INDIA


In 2017, the real GDP per person was 98,028 rupees.

In 2018, the real GDP per person was 104,191 rupees.

Calculate the real GDP per person growth rate -

Growth rate = [(Real GDP per person in 2018 - Real GDP per person in 2017)"\div" Real GDP per person in 2017] "\times" 100

Growth rate "= \frac{104,191 - 98,028}{98,028} \times 100"

Growth rate = 6.29%

India's real GDP per capita is growing at a rate of 6.29 percent.

Calculate the number of years it will take to double real GDP -

Number of years = 70"\div" Real GDP per person growth rate "= \frac{70}{6.29}" = 11.13 years

In around 11 years, India would be able to double its real GDP per person.

It indicates that by 2029, India's real GDP per person will have doubled.

Thus,

India would be the first country to double its level of living if current growth rates are maintained.

By 2029, it will be able to double its quality of life.